OpenStudy (anonymous):

lim x >-inf x^5-5000x^4/2x^5 + 7000 Estimate the limit numerically

zepdrix (zepdrix):

numerically? so bust out your dusty calculator, plug in really large negative values, and try to get an idea of what the limit is approaching.

OpenStudy (anonymous):

Thanks I was not going big enough!!

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} (x ^{5} - x ^{4}5000) / 2x ^{5}+7000\] \[=\lim_{x \rightarrow -\infty} (( x ^{5}) (1-5000/x))/ x ^{5} (1+7000/x ^{5})\]

OpenStudy (anonymous):

cancel out x^5 and plug x = - infty your answer is approx 1/2

OpenStudy (anonymous):

I got 2.1 does that sound right??

OpenStudy (anonymous):

0.5